OpenStudy (anonymous):

A sound wave of intensity I>1 has an amplitude 30 mm. Another wave of the same frequency but of intensity I>2, has an amplitude of I 0 mrn. What is the value of (I>1/I>2) ? (A) 3 (B) 9 (C) 27 (D) 81

OpenStudy (anonymous):

P = 0.5(omega)^2A^2(rho)v = 0.5(2*pi*f)^2*A^2(rho)v so P is variable to A^2 since v,f,and others are constant I>1/I>2 = P1/P2 = (30/10)^2 = 9

OpenStudy (anonymous):

thank u